OpenStudy (dmezzullo):

2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z

OpenStudy (anonymous):

Is this supposed to be an equation? Or are you supposed to simplify it.

OpenStudy (swissgirl):

All you gotta do is gather like terms 2x-2x-2y-y+5z+3z

OpenStudy (dmezzullo):

combine like terms

OpenStudy (anonymous):

huh?

OpenStudy (dmezzullo):

c?

OpenStudy (swissgirl):

Close its the 4th option 2x-2x=0 -2y-y=-3y 3z+5z=8z -3y+8z is your answer

OpenStudy (dmezzullo):

thanks